# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

### v2

@@ IV.1 Dwell time
$T_{D}(\tau)=\frac{1}{2\pi}\int_{0}^{\tau}\theta(x(t))dt\ .$ (4.1)
The dwell time is a measure of the fraction of time the particle spends in $x>0$. Classically, this value should be $1/2$ over a full period ($\tau=2\pi$). This can readily be turned into a self-adjoint operator, $\hat{T}_{D}(\tau)$ [73]. The average quantum-mechanical dwell time $\langle\hat{T}_{D}(\tau)\rangle$ can be ascertained experimentally by measuring $\langle\hat{\theta}(x(t))\rangle$ at multiple times.
@@
$\langle\hat{T}_{D}(2\pi)\rangle=\sum_{n}|c_{n}|^{2}\langle n|\hat{\theta}(t)|n\rangle\ .$ (4.2)
As $|n\rangle$ are stationary states, $\langle n|\hat{\theta}(t)|n\rangle=\langle n|\hat{\theta}(0)|n\rangle$, and because energy eigenstates have definite parity, $\langle n|\hat{\theta}(0)|n\rangle=\frac{1}{2}$. Therefore,
$\langle\hat{T}_{D}(2\pi)\rangle=\frac{1}{2}.$ (4.3)
@@
[…]
It is quite striking that the quantum dwell time agrees with the classical result and yet Tsirelson violations still occur. In fact, Ref. [4] considered alternatives to the Tsirelson scheme with a greater number of measurement times during a period and found that quantum violations faded away as the number of measurement times was increased (limiting to the dwell time result which has no quantum violations). This perspective motivates viewing the Tsirelson quantity as a Riemann sum approximation to the dwell time integral. Upon performing such an analysis (see Appendix K), we have found that the Tsirelson quantity is related to the probability density current, $J(x,t)=\operatorname{Im}\left(\psi^{*}\partial_{x}\psi\right)$, at the origin:
$\langle A\rangle=(6\langle\hat{T}_{D}(2\pi)\rangle-3)+2\pi J(0,\tau)=2\pi J(0,\tau)\ ,$ (4.4)
where $\tau$ is some time on the interval $0,2\pi]$, as the result was obtained using the mean value theorem. Thus, we see that the dependence of $\langle A\rangle$ upon the average dwell time drops out, since $\langle\hat{T}(2\pi)\rangle$ assumes the classical value, and we see that the Tsirelson violation is fundamentally rooted in the larger-than-classical probability current at the origin.
IV.2 Crossing number
@@
Motivated by this classical expression, we define the Hermitian quantum operator
@@
where $\operatorname{sign}(\hat{p}(t))$ and $\delta(\hat{x}(t))$ are defined via the functions being applied to the eigenvalues of the spectral decomposition. We can expand out $\hat{N}_{c}(\tau)$ as
$\hat{N}_{c}(\tau)=\frac{1}{4}\int_{0}^{\tau}\mathrm{d}t\iint\mathrm{d}p\ \mathrm{d}q\operatorname{sign}(p)qe^{iHt}\left[A+B\right]e^{-iHt}\ ,$ (4.7)
where
$A=|p\rangle\langle p|q\rangle\langle q|x_{0}\rangle\langle x_{0}|+|x_{0}\rangle\langle x_{0}|q\rangle\langle q|p\rangle\langle p|\ ,$ (4.8)
$B=|q\rangle\langle q|x_{0}\rangle\langle x_{0}|p\rangle\langle p|+|p\rangle\langle p|x_{0}\rangle\langle x_{0}|q\rangle\langle q|,$ (4.9)
and where $\langle x|p\rangle=\frac{1}{\sqrt{2\pi}}e^{ipx}$ and $|x_{0}\rangle$ is an eigenstate of the position operator at the origin. Then, when $\tau=2\pi$, the operator is diagonal in an energy eigenbasis and the diagonal elements are
$\langle n|\hat{N}_{c}(2\pi)|n\rangle=\frac{\pi}{2}\iint\mathrm{d}p\ \mathrm{d}q\operatorname{sign}(p)q\left[\langle n|A|n\rangle+\langle n|B|n\rangle\right]\ .$ (4.10)
But more conveniently, when Eq. 4.6 is written as
[…]
we note that the first term picks up only symmetric states and the second term only anti-symmetric states.
Explicit calculations indicate an asymptotic tendency towards the classical value of two crossings in the limit of high energy eigenstates, which is depicted in Fig. 6. Moreover, we calculate for the Zaw–Scarani state an average crossing number of $\langle\hat{N}_{c}(2\pi)\rangle_{Z}\approx 1.49$, with a standard deviation of $\Delta\hat{N}_{c}(2\pi)_{Z}\approx 0.2$. Over a third of a period, between measurement times, $\hat{N}_{c}$ remains diagonal for states in $\mathcal{H}_{0}$, and so $\langle\hat{N}_{c}(2\pi/3)\rangle_{Z}\approx 0.50$ with a standard deviation of $\Delta\hat{N}_{c}(2\pi/3)_{Z}\approx 0.07$. These results indicate that the probabilities of non-classical behaviour, such as two crossings during a third of a period, are very small indeed.
Figure 6: Plot of the average crossing number, $\langle n|\hat{N}_{c}(2\pi)|n\rangle$ (see Eq. 4.10), over a full period for the first 21 energy eigenstates. We see that the even eigenstates converge much more rapidly to the classical value of $2$.
@@
in a fixed energy state. This in turn means that a semi-classical approximation can be made for symmetric states, wherein the first term of Eq. 4.11 can be taken to be $\sim\sqrt{2E}|\psi(0)|^{2}$. Via this approximation, only single-time measurements are required. However, states like the Zaw–Scarani contain a large anti-symmetric part; also, the spread $\Delta\hat{N}_{c}$ cannot be determined with single-time measurements, and thus this approach is only partially successful.
IV.3 Implications for trajectory models
@@
We have seen that the quantum-mechanical properties of the dwell time are identical to its classical properties and that as for the crossing number, its quantum-mechanical properties are qualitatively the same as classical and quantitatively similar. These features suggest that measurements of the dwell time and crossing number may be used to give indirect but useful bounds on the size of uniform precession violations. These quantities are, in principle, related to the three-time probabilities used in Section III, but any such relation will be non-unique, and we have not been able to find a particularly useful one. We proceed using a more heuristic analysis.
@@
We will suppose that there exists an underlying probability distribution on a set of phase-space trajectories $p(t)$, $x(t)$ which are close to classical trajectories but could potentially violate the Tsirelson inequality by violating uniform precession. We then examine what restrictions on these trajectories are implied by measurements of the dwell time and crossing number, with a view to ruling out these alternative classical explanations of the Tsirelson violations.
The properties of the dwell time, Eq. 4.3, over the time interval $0,2\pi]$ indicate that the underlying trajectories must each spend equal amounts of time in $x>0$ and $x<0$. However, as noted, the dwell time must have the same properties over any shifted time interval of the same size. The implication in phase space is the trajectories must spend equal times above or below any infinite radial line. Thus, the set of trajectories must have rotational symmetry, which suggests that the only possibility is a set of circles around the origin with constant angular velocity along each circle. This still allows a set of circular trajectories with a spread of possible constant angular velocities, which will cause a violation of uniform precession. However, trajectories with angular velocities faster or slower than the classical one for the given Hamiltonian will create non-classical values for the crossing number. But, as we have seen above, the probabilities for such non-classical values of crossing numbers are very small.
IV.4 Classical versus quantum dynamics
Finally, we make some general remarks about dynamics in classical versus quantum theories and why we might expect, on general grounds, that the violation of uniform precession be small. The predictions of quantum theory are very different to the predictions of classical physics, and most of these differences trace back to the simple fact that quantum theory allows a much richer variety of initial states than does classical physics—for example, through Wigner functions that can be negative versus classical phase-space probabilities. This fact is responsible for some of the most striking quantum phenomena, such as Bell inequality violations.
However, classical and quantum physics look at lot more similar at the level of dynamics, and there are many results in the history of quantum theory that prove this. For example, wave packets—or more generally, phase-space localised quasi-projectors—evolve to a good approximation along classical paths (see for example Ref. [74] and references therein). This in turn is related to the fact that the Wigner function, for a class of potentials, has approximately classical evolution. This behaviour can be affected by monitoring (e.g. with projection operators onto ranges of position at intermittent times), but this effect is not large unless the projections are onto very small regions and the projectors are applied at time intervals shorter than the energy time $\hbar/\langle E\rangle$ [75]. (For the Zaw–Scarani state considered in this paper, the energy time is calculated to be $1/12$ of a period.) Furthermore, in specific models, such as LG2 violations for a Gaussian intial state with harmonic oscillator dynamics, there is some non-classicality, but only about $30\%$ of the level of non-classicality that can be produced by a negative Wigner function [65, 66]. Of course, this discussion invokes quantum-mechanical properties, but the sceptical macrorealist could confirm these ideas using control experiments in which the dynamics of a system in a set of initially classical-like states are measured.
@@ IV.4 Classical versus quantum dynamics
So, in summary, there are general physical and experimentally-testable reasons why we might expect the violation of a dynamical assumption like uniform precession to be reasonably small compared to the more significant non-classical effects that can arise from the initial state.
V Summary and Discussion
It is remarkable that a simple assumption about the classical dynamics of an oscillator—so straightforward that one would think it must surely be upheld universally—in fact reveals ever more of the bizarre and distinguishing predictions of quantum theory. The Tsirelson inequality also makes clear that, by introducing a dynamical assumption, the prospect of observing non-classicality in temporal correlations need not be plagued by the problem of circumventing invasive sequential measurements.
@@ V Summary and Discussion
In Section II, we analysed maximally-violating states and formalised the Tsirelson operator as the twirl of the sign operator over the group elements of the three-element cyclic group. This perspective directly explained its eigenspace structure and opens up the possibility for more tools from quantum information to provide further insight into the Tsirelson operator. We provided proofs that a superposition of at least three energy eigenstates of the QHO is required to produce a Tsirelson violation and that, in relevant subspaces, all three correlators are equal for the specific measurement time interval of two-thirds of a period. The violations possible with simple states such as the Zaw-Scarani state—and, moreover, superpositions of coherent states—are certainly promising and within the reach of experimentalists in the coming years.
@@
In Section III, we broke apart the Tsirelson quantity into separate terms that quantified the breaking of uniform precession and quantum interferences, such that their sizes may be compared. Recognising that uniform precession can be defined quantum-mechanically in several inequivalent ways in terms of measurable probabilities, we investigated two methods. In the first method, we proved the surprising fact that eigenstates of the Tsirelson operator satisfy all four LG3 inequalities, meaning that there invariably exists a macrorealistic description of sign changes, which can then be used to assess uniform precession through an LG3 test. In the second method, we considered a generalised dynamical assumption with the difference of no-sign-change probabilities (perhaps less desirable than the first method, for it involves three sequential measurements). In all cases, we found success with the protocol, demonstrating that quantum interference terms must be implied, elevating the Tsirelson test from a simple test of non-classical behaviour to a more robust demonstration of quantumness.
In Section IV, by considering the heuristic notions of dwell times and the crossing number operator, we presented further qualitative evidence that uniform precession is satisfied well in states that violate the Tsirelson inequality. We found that the quantum-mechanical dwell time obeys the classical bound exactly and that the total number of origin crossings is approximately the classical value. On their own, these stand as interesting mathematical results, further elucidating the various ways in which uniform precession is and is not equivalent in the classical and quantum cases. But we have also further interpreted these quantities in the context of underlying trajectory models, noting that they greatly restrict the extent to which underlying trajectories can be erratic. We also came upon the global property that Tsirelson violations are linked to the probability currents through the origin, whereby weakly-flowing states, or the classicalised currents of quantum states, never produce a violation.
@@
However, an ultimate goal for Tsirelson tests will be to probe systems not knowing beforehand whether they are quantum or not. Then, to an experimentalist—who should wish to implement this protocol by collecting a data set of the Tsirelson quantity supplemented with a Leggett–Garg test for the relevant probabilities—the problem of re-invoking a Leggett–Garg test, in order that we should further conclude that macrorealism per se has been violated, is possibly harder to justify. But we may appeal here to the heuristic results of Section IV, which provide evidence that the LG3 quantity $L_{1}$ should be small and which support the idea that violations of uniform precession are not, in general, sufficiently large to be the primary cause of Tsirelson inequality violations. In some cases, we were partially successful in showing how these quantities may be measured through only single-time measurements. In the original Leggett–Garg scheme, there is no reason to expect a priori that $L_{1}$ should be small; here, however, we have gathered evidence that it should indeed be small (and positive), and so if an experiment finds otherwise, there is reason to suspect that measurement invasiveness may be largely to blame.
The results of this paper open up several new research directions. Future work could consider an anharmonic oscillator, multiple coupled oscillators and links to field theories, entanglement, alternative measures of crossings, and stronger quantum-mechanical definitions of uniform precession, amongst other things. We have concentrated in this work upon the case of the harmonic oscillator, but it would be interesting to conduct this same analysis and assessment of uniform precession for spin systems; subsequent comparison with the results herein obtained could be insightful. The idealistic and experimentally impracticable nature of perfect sign operator projective measurements could be remedied by extending this analysis to use smoothed projectors. The analysis of Tsirelson’s inequality and the breaking of uniform precession could be extended to incorporate suspected connections with the quantum back-flow limit. This could all provide further illumination on the true nature of violations of Tsirelson’s intriguing inequality and, altogether, bring us closer to loophole-free tests of quantum behaviour.

### v3

@@ V.1 Dwell time
$T_{D}(\tau)=\frac{1}{2\pi}\int_{0}^{\tau}\theta(x(t))dt\ ,$ (5.1)
where we highlight the explicit dependence of $\theta$ upon the time-dependent position $x(t)$. The dwell time is a measure of the fraction of time the particle spends in $x>0$. Classically, this value should be $1/2$ over a full period ($\tau=2\pi$). This can readily be turned into a self-adjoint operator, $\hat{T}_{D}(\tau)$ [Munoz2009DwellTime]. The average quantum-mechanical dwell time $\langle T_{D}(\tau)\rangle$ can be ascertained experimentally by measuring $\langle\theta(\hat{x}(t))\rangle$ at multiple times.
@@
$\langle T_{D}(2\pi)\rangle=\sum_{n}|c_{n}|^{2}\langle n|\theta(\hat{x}(0))|n\rangle\ ,$ (5.2)
because $|n\rangle$ are stationary states, meaning $\langle n|\theta(\hat{x}(t))|n\rangle=\langle n|\theta(\hat{x}(0))|n\rangle$. Also, as energy eigenstates have definite parity, $\langle n|\theta(\hat{x}(0))|n\rangle=\frac{1}{2}$. Therefore,
$\langle T_{D}(2\pi)\rangle=\frac{1}{2}.$ (5.3)
@@
The dwell time operator is diagonal over a full period (and in fact is proportional to the identity)—which can be understood as due to the fact that $\hat{T}_{D}(2\pi)$ is $)\text{Twirl}(\theta(\hat{x}$)) over a representation of the group U(1)U(1)—and thus there is no spread in its distribution. Within the subspace $\mathcal{H}_{0}$, the dwell time operator is also diagonal when integrated to one-third or two-thirds of a period. The latter case is particularly informative, since this represents integrating between the first and final measurement times. We have found that $\langle\hat{T}_{D}(4\pi/3)\rangle=\frac{1}{3}$ exactly—which, remarkably, also falls within the range of classically allowed values. Another observation is that looking at the dwell time over any cycle, starting at any time, yields the same results.
It is quite striking that the quantum dwell time agrees with the classical result and yet Tsirelson violations still occur. In fact, Ref. [precessions] considered alternatives to the Tsirelson scheme with a greater number of measurement times during a period and found that quantum violations faded away as the number of measurement times was increased (limiting to the dwell time result which has no quantum violations). This perspective motivates viewing the Tsirelson quantity as a Riemann sum approximation to the dwell time integral. Upon performing such an analysis (see Appendix L), we have found that the Tsirelson quantity is related to the probability density current, $J(x,t)=\operatorname{Im}\left(\psi^{*}\partial_{x}\psi\right)$, at the origin:
$\langle A\rangle=(6\langle T_{D}(2\pi)\rangle-3)+2\pi J(0,\tau)=2\pi J(0,\tau)\ ,$ (5.4)
where $\tau$ is some time on the interval $0,2\pi]$, as the result was obtained using the mean value theorem. Thus, we see that the dependence of $\langle A\rangle$ upon the average dwell time drops out, since $\langle T(2\pi)\rangle$ assumes the classical value, and we see that a Tsirelson violation necessitates a ‘large’ (i.e. $1/2\pi$) probability current at the origin. Consequently, this relation restricts the choice of violating states. Conversely, Eq. 5.4 provides a condition guaranteeing Tsirelson satisfaction. Specifically, if $|J(0,\tau)|\leq 1/2\pi$ for all $\tau\in[0,2\pi]$, then $|\langle A\rangle|\leq 1$. In Appendix L, we also consider classicalisations of the quantum phase-space distribution and the resulting classical probability current. These classicalised currents were consistently lower than the corresponding quantum currents, indicating a connection to the above condition on Tsirelson satisfaction.
V.2 Crossing number
@@
Likewise, we can define the Hermitian quantum operator
@@
suppressing the time dependence of $\hat{x},\hat{p}$. We note that, written in this form, the first term picks up only symmetric states and the second term only anti-symmetric states when $\tau=2\pi$.
Explicit calculations indicate an asymptotic tendency towards the classical value of two crossings in the limit of high energy eigenstates, which is depicted in Fig. 6. Moreover, we calculate for the Zaw–Scarani state an average crossing number of $\langle N_{c}(2\pi)\rangle_{Z}\approx 1.49$, with the possible measurement outcomes $\langle 0|\hat{N}_{c}(2\pi)|0\rangle=\langle 3|\hat{N}_{c}(2\pi)|3\rangle=\sqrt{2}$ and $\langle 6|\hat{N}_{c}(2\pi)|6\rangle\simeq 2.033$ (as the energy eigenstates are also eigenstates of $\hat{N}_{c}(2\pi)$). So, the spread of possible measurement outcomes is around $0.6$. Over a third of a period, between measurement times, $\langle N_{c}(2\pi/3)\rangle_{Z}\approx 0.50$. These results indicate that the number of crossings in the quantum case for the Zaw–Scarani state is significantly smaller than the classical case.
Figure 6: Plot of the average crossing number, $\langle n|N_{c}(2\pi)|n\rangle$, over a full period for the first 21 energy eigenstates. We see that the even eigenstates converge much more rapidly to the classical value of $2$.
@@
in a fixed energy state. In turn, the first term of Eq. 5.6 can be taken to be proportional to $\sqrt{2E}|\psi(0)|^{2}$, where we include the non-zero ground state energy contribution in $E$. This provides an alternative expression for the crossing number for fixed energy states which only involves a single-time measurement. Using this form, we find that $\langle 0|\hat{N}_{c}|0\rangle\approx 1.7725$ and $\langle 6|\hat{N}_{c}|6\rangle\approx 1.997$. These results are reasonably close to those obtained above and exhibit the same qualitative features. However, this method only works for the symmetric states, and we note that states like the Zaw–Scarani state contain a large anti-symmetric part.
V.3 Implications for trajectory models
@@
We have seen that the quantum-mechanical properties of the dwell time are identical to its classical properties and that as for the crossing number, its quantum-mechanical properties are qualitatively the same as classical and quantitatively similar. These features suggest that measurements of the dwell time and crossing number may be used to give indirect but useful bounds on the size of uniform precession violations. These quantities are, in principle, related to the three-time probabilities used in Section IV, but any such relation will be non-unique, and we have not been able to find a particularly useful one. We proceed using a more heuristic analysis.
@@
We will suppose that there exists an underlying probability distribution on a set of configuration space trajectories $x(t)$ which are close to classical trajectories but could potentially produce a Tsirelson inequality violation by violating uniform precession. A trajectory model could produce non-zero values of $p(+,+,+)$ and p(,,)p(-,-,-) if some fraction of the trajectories are on the same side of the $x$-axis at all three measurement times. We argue that the dwell time and crossing number results limit such alternative explanations.
The properties of the dwell time, which hold for any time interval covering a single period and for any initial state, indicate that the underlying trajectories must spend equal times in $x>0$ and $x<0$. Hence, over two thirds of a period, each trajectory cannot spend more than half a period on one side, so must cross at least once between the first and third measurement times. This means that there can be no trajectories remaining on one side for two thirds of a period. So, the only trajectories contributing to $p(+,+,+)$ and p(,,)p(-,-,-) must, in order to be on the same side at all three measurement times, cross at least twice during the first third of a period, or during the second third, or both. To rule out such wiggly trajectories, we can appeal to the crossing operator results which indicate that highly erratic trajectories have very low probability.
This argument gives an independent understanding as to why uniform precession should hold reasonably well. We have not found a more precise quantitative account to compare with the results of Table 2.
V.4 Classical versus quantum dynamics
Finally, we make some general remarks about dynamics in classical versus quantum theories and why we might expect, on general grounds, that the violation of uniform precession is small.
The predictions of quantum theory are very different to the predictions of classical physics and there are a number of reasons for this. Quantum theory allows a much richer variety of initial states than does classical physics—for example, through Wigner functions that can be negative versus classical phase–space probabilities. In general, there is also a wide variety of possible single-time measurements that can be made that contribute to quantum behaviour. For two-time sequential measurements, the non-commutativity of measurements also contributes.
However, when assessing the degree to which the dynamics of a quantum-mechanical point-particle system is behaving classically, we are only interested in a very limited class of measurements, namely, coarse-grained position or phase–space measurements at different times. For such measurements, the dynamics of classical and quantum systems actually look a lot more similar. There are standard results showing that wave packets and, more generally, phase–space localised quasi-projectors evolve to a good approximation along classical paths (see for example Ref. [Omnes1997] and references therein). These results can be used to show that the probabilities for histories of coarse-grained position or phase–space measurements are strongly peaked around classical paths.
This near-classical behaviour is modified if the measurements significantly chop the evolving wave function, as is the case in this paper (and, for multiple measurements, if they are applied at time intervals shorter than the energy time $\hbar/\langle E\rangle$ [Omnes1997], but this is not significant here). The magnitude of the non-classical effects producing the chopping can be seen in the specific model of Refs. [MawbySHO, PhysRevA.107.032216], which exhibited LG2 violations for measurements of $\text{sign}(x)$ for a harmonic oscillator with Gaussian initial state. This was produced by the LG2 corresponding to the quasi-probability q(,+)q(-,+), which we write explicitly in the Wigner representation:
$q(-,+)=\int dp\ dq\ W_{-+}(p,q)W_{\rho}(p,q)\ ,$ (5.8)
where $W_{\rho}$ is the Wigner function of the Gaussian initial state, which is clearly non-negative, and $W_{-+}$ is the Wigner–Weyl transform of the operator $(1/2)\{\theta(\hat{x}(t_{1})),\theta(\hat{x}(t_{2})\}$, given in Ref. [PhysRevA.107.032216]. The classical analogue of $W_{-+}(p,q)$ is clearly $\theta(x(t_{1})\theta(x(t_{2}))$, which is trivially non-negative; but $W_{-+}(p,q)$ has regions of negativity which arise from the non-commutativity of the two operators, and this is the only source of non-classicality. (Physically, this is related to the diffraction-in-time effect [PhysRevA.107.032216, PhysRev.88.625, moshinsky_diffraction_1976].) Most significantly, the violation produced is only $22\%$ of the maximum possible violation (the Lüders bound). By contrast, near-maximal LG violations are possible in this system with, for example, an initial state consisting of the first excited state of the harmonic oscillator, which has a Wigner function with substantial regions of negativity. (More speculatively, we also note that the size of the uniform precession violation here is roughly $4\%$, which is peculiarly similar to the estimated bound on the so-called quantum back-flow effect [bracken_probability_1994, halliwell_quantum_2013, PhysRevA.99.012124, trillo_quantum_advantage].)
This example indicates why we might expect that, for the very limited set of measurements used to check proximity to classical behaviour, the non-classical effects arising from those measurements will generally not match the level of non-classicality arising from the choice of initial state. These assertions could be checked experimentally using control experiments in which the dynamics of a system prepared in a set of classical-like states are measured.
@@ V.4 Classical versus quantum dynamics
So, in summary, there are general physical and experimentally-testable reasons why we might expect the violation of a dynamical assumption like uniform precession to be small compared to the more significant non-classical effects that can arise from the initial state.
VI Summary and Discussion
The Tsirelson inequality makes clear that, by introducing a classically-motivated dynamical assumption, the prospect of detecting quantum behaviour in temporal correlations need not be plagued by potentially invasive sequential measurements. In this paper, we have placed the Tsirelon inequality within the wider framework of Leggett-Garg tests for macrorealism, thereby yielding a method for an assessment of the dynamical assumption.
@@ VI Summary and Discussion
In Section III, we analysed violating states in an energy-truncated subspace and formalised the Tsirelson operator as the twirl of the sign operator over the group elements of the three-element cyclic group. This perspective directly explained its eigenspace structure and opens up the possibility for more tools from quantum information to provide further insight into the Tsirelson operator. We provided proofs that a superposition of at least three energy eigenstates of the QHO is required to produce a Tsirelson violation and that, in relevant subspaces, all three correlators are equal for the specific measurement time interval of two-thirds of a period. The violations possible with simple states such as the Zaw-Scarani state—and, moreover, superpositions of coherent states—are certainly promising and within the reach of experimentalists in the coming years.
@@
In Section IV, we broke apart the Tsirelson quantity into separate terms that quantified the breaking of uniform precession and quantum interferences, such that their sizes may be compared. Recognising that uniform precession can be defined quantum-mechanically in several inequivalent ways in terms of measurable probabilities, we investigated two methods. In the first method, we proved the surprising fact that violating eigenstates of the Tsirelson operator satisfy all four LG3 inequalities, meaning that there invariably exists a macrorealistic description of sign changes, which can then be used to assess uniform precession through an LG3 test. In the second method, we considered a generalised dynamical assumption with the difference of no-sign-change probabilities (perhaps less desirable than the first method, for it involves three sequential measurements). In all cases, we found success with the protocol, demonstrating that quantum interference terms must be implied, elevating the Tsirelson test from a simple test of non-classical behaviour to a more robust demonstration of quantumness.
In Section V, by considering the heuristic notions of dwell times and the crossing number operator, we presented further qualitative evidence that uniform precession is satisfied well in states that violate the Tsirelson inequality. We found that the quantum-mechanical dwell time obeys the classical bound exactly and that the total number of origin crossings is approximately the classical value. On their own, these stand as interesting mathematical results, further elucidating the various ways in which uniform precession is and is not equivalent in the classical and quantum cases. But we have also further interpreted these quantities in the context of underlying trajectory models, noting that they greatly restrict the extent to which underlying trajectories can be erratic. We also came upon the global property that Tsirelson violations are linked to the probability currents through the origin, whereby weakly-flowing states, or the classicalised currents of quantum states, never produce a violation.
@@
However, an ultimate goal for Tsirelson tests will be to probe systems not knowing beforehand whether they are quantum or not. Then, to an experimentalist—who should wish to implement this protocol by collecting a data set of the Tsirelson quantity supplemented with a Leggett–Garg test for the relevant probabilities—the problem of re-invoking a Leggett–Garg test, in order that we should further conclude that macrorealism per se has been violated, is possibly harder to justify. But we may appeal here to the heuristic results of Section V, which provide evidence that the LG3 quantity $L_{1}$ should be small and which support the idea that violations of uniform precession are not, in general, sufficiently large to be the primary cause of Tsirelson inequality violations. In some cases, we were partially successful in showing how these quantities may be measured through only single-time measurements. In the original Leggett–Garg scheme, there is no reason to expect a priori that $L_{1}$ should be small; here, however, we have gathered evidence that it should indeed be small (and positive), and so if an experiment finds otherwise, there is reason to suspect that measurement invasiveness may be largely to blame.
The results of this paper open up several new research directions. Future work could consider an anharmonic oscillator, multiple coupled oscillators and links to field theories, entanglement, alternative measures of crossings, and stronger quantum-mechanical definitions of uniform precession, amongst other things. We have concentrated in this work upon the case of the harmonic oscillator, but it would be interesting to conduct this same analysis and assessment of uniform precession for spin systems; subsequent comparison with the results herein obtained could be insightful. We have considered the use smoothed projectors, but these order of magnitude arguments need to be bolstered by a more detailed calculation, similar to that of Ref. [MawbySHO]. The analysis of Tsirelson’s inequality and the breaking of uniform precession could be extended to incorporate suspected connections with the quantum back-flow limit. This could all provide further illumination on the true nature of violations of Tsirelson’s intriguing inequality.